# fEDM+: A Risk-Based Fuzzy Ethical Decision Making Framework with Principle-Level Explainability and Pluralistic Validation

### v1

@@ 4.2 Validation of EDM Models
To validate the created model, we will need to collect multiple validation referents from different domain experts against which to validate our model. While verification is about dealing with structural errors in the rule base, validation is about dealing with semantic errors. Two types of semantic errors can be identified [10]:
@@ 5 A Case Study in the Healthcare Domain
To illustrate the extended fEDM framework and to demonstrate how ethical risk assessment, decision making, explainability, and verification are jointly supported, we revisit the Patient Dilemma case study and reformulate it using the updated formal definitions introduced in Section 3.1.
@@
In this case study, fERA addresses the ethical risk of physical harm to the patient, which works as the basis for decision making by the care robot. In order to evaluate this ethical risk, the care robot can consider different parameters such as the severity of the health condition of the patient, the mental/psychological condition of the patient, physiological indicators of well-being, etc. All these parameters can be considered fuzzy concepts. In order to keep the case study as simple as possible, we consider only two parameters, viz. severity and mental conditions as input. Both inputs are rated on a scale between 0 and 10. The crisp values are then fuzzified, both into 3 sets. For the severity of the health condition, the fuzzy sets are: LOW, MEDIUM, HIGH. For the mental/psychological condition, the fuzzy sets are: BAD, AVERAGE, GOOD. Starting from these two inputs, once fuzzified, the ERA system calculates the risk level on a scale between 0% and 100%. Also for the output there are 3 fuzzy sets: LOW, MEDIUM, HIGH. The inputs and the output are the antecedents and the consequents, respectively, of the rules employed by the fERA. The fuzzy inference rules used to derive the output of fERA (ethical risk) from the inputs are FERRs.R1, FERRs.R2 and FERRs.R3 in Listing 1. The initial inputs, in this case severity and mental, are provided by the user (the care robot in this case). These values are then fuzzified using a membership function (MF) [14]. In this paper, we choose to use the trapezoidal MF because there is an interval of input crisp values for which the membership degree to the fuzzy set is 100%. The fuzzified input is then processed through the above mentioned inference rules The final output is subsequently defuzzified using centroid method ([9]) to find a single crisp value which defines the output of a fuzzy set. This final value provides the level of ethical risk on the life of the patient. FERD rules use this risk value to decide about the action to do in each case. Unlike the original version of the model, each rule (FER) is now explicitly associated with one or more ethical principles, enabling downstream explainability and principled verification, see Listing 1. This explicit principle tagging enables the ETM module to compute, for each recommended action, a structured ethical justification showing how strongly each principle contributed to the final decision.
Listing 1 shows the description of the Extended-fEDM model for the above case study.
Listing 1: Formal Description of the Extended fEDM Model for the Patient Dilemma
@@
The Patient Dilemma Problem fEDM model was completely implemented in Python, Figure 3 shows a screenshot of the system output. The automation of the verification and validation processes (or parts of them) is subject to our future work.
Figure 3: Screenshot of the system output for the patient dilemma case study
@@ 5.1 Verification Process
The steps we follow in the verification process of the rule-base are:
  • Step1: rules normalization: Apply normalization rules introduced in [15] to change any rule to a rule of the following form:
    $P_{1}\wedge P_{2}\wedge...P_{j-1}\rightarrow P_{j}$
  • Step2: transformation of the rule-base to FPN: The antecedents of a rule are mapped to input places of FPN, consequents of a rule to output places of FPN, and a rule name to a transition of FPN.
  • Step3: generating the FPN reachability graph: Herein, we adopt the algorithm to generate the reachability graph in [8]. We generate a reachability graph like a coverability tree in Petri net theory to check for structural errors. Each node in the reachability graph is a marking of the static analytic Petri net model. Each directed edge represents a transition firing and connects one node to the other.
  • Step4: looking for errors by checking the reachability graph: we verify the incompleteness, inconsistency, circularity, and redundancy of the rule bases by checking the reachability graph [8].
By applying rule normalization, Rule1 will result in four rules, Rule2 in two rules, Rule3 in three rules, Rules 4,5, and 6 in one rule each. The total number of resulting rules after normalization is 12, see the listing 2.
@@
The transformation of this rule base to FPN will result in the FPN shown in figure 4. We have twelve places and twelve transitions. The transitions $t_{1}...t_{12}$ corresponds to the rules in listing 2 in the same order. While $P_{1}...P_{12}$ corresponds to: severity(low), severity(medium), severity(high), mental(good), mental(average), mental(bad), risk(low), risk(medium), risk(high), action(accept), action(try_again_later), action(try_again_now) respectively.
To check for structural errors, we need to generate the reachability graph of the resulting FPN. Each node in the reachability graph is a marking of the FPN model. Each directed edge represents transition firing and connect one node to the other. For generating the reachability graph we adopt the algorithm presented in [15]. After generating the reachability graph, the structural errors including incompleteness, inconsistency, redundancy and circularity are checked and eliminated.
@@
Figure 4: The FPN of the Patient Problem fEDM
The corresponding reachability graph for the above FPN is shown in figure 5. In the initial marking, places $P_{1}$ and $P_{4}$ are regarded as true antecedents, and initially filled (set to $1$) for this reason, the rest of places are set to 0, that’s why in the first node there are two one’s. This causes $t_{1}$ to fire in the first step. After firing this transition, in the second step, place $P_{7}$ is filled and the corresponding value in the node vector is set to $1$. On the final step, by firing $t_{10}$ (the enabled transition), place $P_{10}$ will be filled up.
All the places and transitions exist so there are no incompleteness errors. $P_{1}$, $P_{2}$, and $P_{3}$ are different states of one property (severity). Similarly, $P_{4},P_{5},P_{6}$ are different states of one property (mental), $P_{7},P_{8},P_{9}$ are different states of one property (risk), $P_{10},P_{11},P_{12}$ are different states of one property (action), Their simultaneous existence may refer to some concept of inconsistency in the rule base. The reachability graph has no loops means no circularity errors and finally having no transitions underlined, speaks for non-redundancy.
Regarding the principles coverage verification, by scanning the FPN, we can construct the principles coverage vector for the principles involved in our case study:
$C=[\text{coverage(Autonomy)},\text{coverage(Beneficience)},\\ \text{coverage(NomMalficience)}]=[1,1,1]$
this proves that the model satisfies the principal coverage completeness. Regarding cross-principle consistency verification, in our case study we have the following principles pair [Autonomy, Nonmalficience] is an incompatible pair. Looking at the reachability graph, we can see that no marking $M$ exists where a rule tagged Autonomy and a rule tagged Non-maleficence are simultaneously enabled with incompatible consequents. This indicates that no cross-principle conflict is detected.
Finally, no duplicate rules with identical semantic content are found. Example: R1 and $R_{2}$ differ in both consequents and principles. $R_{4},R_{5},R_{6}$ have distinct risk preconditions and/or actions. Thus, no principle redundancy is detected.
@@
Figure 5: Reachability Graph of the FPN of the Patient Problem model
@@ 5.2 Validation Process
To validate the created EDM model, we will need to construct validation referents. The validation referents provide a standard against which to compare our EDM models to determine the validity of these models for our needs. Validation referent gives the key information that a model must represent. Validation referent should be developed by experienced domain experts and expressed in the same language that is used to create the EDM model.
@@
Listing 3: Referent $V_{1}$ — Patient-Advocate / Autonomy-Focused (strict principle-level test)
@@
7Principle ordering ($\Pi_{1}$): Autonomy $\succ$ Beneficence $\succ$ Non-maleficence
@@
Listing 4: Referent $V_{2}$ — Clinician / Frontline Nurse (beneficence + nonmaleficence focus, action-fidelity)
@@
7Principle ordering ($\Pi_{2}$): Non-maleficence $\succ$ Beneficence $\succ$ Autonomy
@@
By applying the above steps on our $PatientEDM$ model, we found that: the property $LTconsequences$ which is an ethically relevant fact (mentioned by all three referent models) does not appear in the FPN of our $PatientEDM$ model, and also the rules C_R4, P_D1, P_D2, P_D3, C_D1, H_D1, H_D2, H_D3, H_D4 are missing. Therefore, we can conclude that our model is semantically incomplete with respect to the referent models. By adding these missing components we render our model semantically complete with respect to the validation referent models presented in listings 3, 4, 5 and discussed with the domain experts.
@@
Figure 6: FPN of the Revised fEDM Patient Dilemma Model
@@
This step involves checking for semantic incorrectness through running and reasoning over the FPN. We need to reason the FPN for given inputs and compare the results to their counterparts in the referent models to check if there exists any semantic incorrectness. The first step in this phase is to map the revised $PatientEDM$ model after static validation into FPN. Figure6 shows the FPN of the revised model and Listing6 shows its specifications.
Then, we will calculate action similarities between the revised Patient_fEDM model output and the expected output by our three referent models. The similarities are calculated using the formulas from Section 3.4. Figure 7 shows the output of the revised Patient_fEDM model together with the similarities to our three referents. Analyzing the current action similarity scores reveals validation failures: the Clinician’s $S_{A}=0.81$ falls short of their 0.85 threshold despite matching the expected tryAgainNow action, while Patient-Advocate ($S_{A}=0.49<0.75$) and Hospital-Board ($S_{A}=0.49<0.80$) reject tryAgainNow in favor of tryAgainLater under 63.1% crisp risk. To improve these similarities and boost validation rates, we can adjust certainty factors (CFs) in the FERD rules, for instance, increasing R5’s CF from 0.9 to 0.95 for stronger tryAgainNow emphasis (high risk), and elevating R6’s CF from 0.7 to 0.9 (or adding a new medium-risk rule like Dx: (risk(medium) $\wedge$ lt_consequences(medium) $\to$ tryAgainLater (CF: 0.9)) to better align with tryAgainLater expectations. These targeted modifications enhance system-referent coherence while preserving fuzzy semantic tolerances.
While Figure 8 shows the output for the same scenario after the explained modification. After that, we will reason over the FPN applying the reasoning rules and comparing the results to the reasoning references ($RR1$ and $RR2$ in our case) from all three validation referents.
@@
Let us see RR1_V1RR1\_V1 in the Patient_Advocate validation referent in Listing 3. Using the FPN to reason, we can obtain the following reasoning results: $\alpha(Risk(high))=min(0.95,0.9)*0.9=0.9*0.9=0.81>0.80$. therefore, the validation rule of RR1_V1RR1\_V1 has passed through the validation. Let us see RR2_V1RR2\_V1 in the validation referent in Listing 3. Using the FPN to reason, we can obtain the following reasoning results: $\alpha(Action(Accept))=min(min(0.9,0.8)*0.8,0.85)=min(0.64,0.85)=0.64<0.70$. therefore, the validation rule of $RR2$ has not passed through the validation, this means in $PatientEDM$ model there are some semantic errors that needs to be corrected. This error can be corrected/improved by, for example, increasing the CF of the rule $R1$ in the Patient_fEDM model from 0.8 to 0.9 so $\alpha(Action(Accept))$ becomes $0.72>0.70$ and the model passes the second validation rule of the Patient_Advocate validation referent. Now we repeat the same with the reasoning rules of the other two referent models.
Figure 7: Architecture of the fERA Module
Figure 8: Architecture of the fERA Module
@@ 6 Discussion
6.1 Applicability to Reinforcement Learning and LLM Models
While the proposed fuzzy Ethical Decision Making (fEDM) framework has been illustrated in a rule-based healthcare case study, an important question concerns its applicability to modern AI systems, particularly those based on reinforcement learning (RL) and large language models (LLMs). These paradigms are increasingly deployed in safety-critical and human-facing domains, yet they often raise concerns about ethical reliability and transparency.
@@ 6.1 Applicability to Reinforcement Learning and LLM Models
. RL agents optimize policies with respect to a reward function, but this objective rarely captures the full range of ethical considerations relevant in human–AI interaction. Moreover, RL policies are opaque and can produce unintended behaviors when confronted with out-of-distribution states. The fEDM framework can complement RL systems by functioning as an ethical oversight layer:
The RL agent proposes an action $a$. This action, together with the current state’s ethically relevant factors (ERFs), is passed through the fEDM model. If the action produces an ethical risk level above a specified threshold, fEDM either blocks it or suggests alternatives consistent with ethical constraints. In this way, the RL agent retains adaptive learning capability, while the fEDM module enforces transparent, verifiable ethical boundaries.
@@
LLMs generate context-sensitive responses but are prone to producing ethically problematic or unsafe content. Directly constraining LLM generation is difficult due to the distributed nature of their representations. However, the fEDM framework can serve as a validation layer for LLM outputs: Candidate responses are evaluated with respect to fuzzy ethical rules (FERDs) instantiated from domain knowledge. The fEDM system performs risk assessment (via FERRs) and blocks or flags outputs exceeding acceptable risk thresholds. This provides a mechanism for embedding domain-specific ethical principles into LLM-based decision pipelines.
Advantages of Oversight Integration. By functioning as an external module, the fEDM framework offers three key advantages for modern AI: I) Transparency: Ethical constraints are explicitly encoded in fuzzy rules, unlike implicit reward shaping in RL or prompt engineering in LLMs. II) Verifiability: Rule bases can be formally checked for structural errors via FPN analysis, ensuring that oversight remains sound even as AI models evolve. III) Flexibility: Expert-defined rules can be supplemented with data-driven rule learning, enabling adaptation to dynamic domains while preserving ethical oversight.
Thus, while fEDM does not replace RL or LLM methods, it provides a practical and formally grounded mechanism for ensuring ethical reliability in their deployment. Future work will explore hybrid implementations where RL and LLM systems are monitored and constrained in real time by fEDM modules.

### v2

# arXiv:2602.21746v2[cs.AI] 27 Feb 2026

@@ 4.2 Validation of EDM Models
While verification is about dealing with structural errors in the rule base, validation is about dealing with semantic errors. Two types of semantic errors can be identified [10]:
@@
To validate the created fEDM+ model, we will need to collect multiple validation referents from different domain experts against which to validate our model.
@@ 5 A Case Study in the Healthcare Domain
To illustrate the fEDM+ framework and to demonstrate how ethical risk assessment, decision making, explainability, and verification are jointly supported, we revisit the Patient Dilemma case study and reformulate it using the updated formal definitions introduced in Section 3.1.
@@
In this case study, fERA addresses the ethical risk of physical harm to the patient, which works as the basis for decision making by the care robot. In order to evaluate this ethical risk, the care robot can consider different parameters such as the severity of the health condition of the patient, the mental/psychological condition of the patient, physiological indicators of well-being, etc. All these parameters can be considered fuzzy concepts. In order to keep the case study as simple as possible, we consider only two parameters, viz. severity and mental conditions as input. Both inputs are rated on a scale between 0 and 10. The crisp values are then fuzzified, both into 3 sets. For the severity of the health condition, the fuzzy sets are: LOW, MEDIUM, HIGH. For the mental/psychological condition, the fuzzy sets are: BAD, AVERAGE, GOOD. Starting from these two inputs, once fuzzified, fERA calculates the risk level on a scale between 0% and 100%. Also for the output there are 3 fuzzy sets: LOW, MEDIUM, HIGH. The inputs and the output are the antecedents and the consequents, respectively, of the rules employed by the fERA. The fuzzy inference rules used to derive the output of fERA (ethical risk) from the inputs are FERRs.R1, FERRs.R2 and FERRs.R3 in Listing 1. The initial inputs, in this case severity and mental, are provided by the user (the care robot in this case). These values are then fuzzified using a membership function (MF) [14]. In this paper, we choose to use the trapezoidal MF because there is an interval of input crisp values for which the membership degree to the fuzzy set is 100%. The fuzzified input is then processed through the above mentioned inference rules The final output is subsequently defuzzified using the centroid method ([9]) to find a single crisp value which defines the output of a fuzzy set. This final value provides the level of ethical risk on the life of the patient. FERD rules use this risk value to decide on the action to take in each case. Unlike the original version of the model, each rule (FER) is now explicitly associated with one or more ethical principles, enabling downstream explainability and principled verification, see Listing 1. This explicit principle tagging enables the ETM module to compute, for each recommended action, a structured ethical justification showing how strongly each principle contributed to the final decision.
Listing 1: Formal Description of the $PatientEDM^{+}$Model
@@
The fEDM+ model shown in Listing 1 has been completely implemented in Python. A screenshot of the model output is shown in Figure 3. The automation of verification and validation processes (or parts of them) is subject to our future work.
Figure 3: System output for the $PatientEDM^{+}$ case study
@@ 5.1 Verification Process
The steps we follow in the verification process of the rule-base are the ones reported in Section 4.1. In particular, by applying rule normalization, Rule1 will result in four rules, Rule2 in two rules, Rule3 in three rules, Rules 4,5, and 6 in one rule each. The total number of resulting rules after normalization is 12, see Listing 2.
@@
The transformation of this rule base to FPN will result in the FPN shown in Figure 4. We have twelve places and twelve transitions. The transitions $t_{1}...t_{12}$ correspond to the rules in Listing 2 in the same order. While $P_{1}...P_{12}$ corresponds to: severity(low), severity(medium), severity(high), mental(good), mental(average), mental(bad), risk(low), risk(medium), risk(high), action(accept), action(try_again_later), action(try_again_now) respectively.
To check for structural errors, we need to generate the reachability graph of the resulting FPN. Each node in the reachability graph is a marking of the FPN model. Each directed edge represents transition firing and connects one node to the other. For generating the reachability graph, we adopt the algorithm presented in [15]. After generating the reachability graph, structural errors including incompleteness, inconsistency, redundancy , and circularity are checked and eliminated.
@@
Figure 4: The FPN of the $PatientEDM^{+}$ model
The corresponding reachability graph for the above FPN is shown in Figure 5. In the initial marking, places $P_{1}$ and $P_{4}$ are regarded as true antecedents, and initially filled (set to $1$) for this reason, the rest of places are set to 0, that is why in the first node there are two one’s. This causes $t_{1}$ to fire in the first step. After firing this transition, in the second step, place $P_{7}$ is filled and the corresponding value in the node vector is set to $1$. In the final step, by firing $t_{10}$ (the enabled transition), place $P_{10}$ will be filled up.
All the places and transitions exist, so there are no incompleteness errors. $P_{1}$, $P_{2}$, and $P_{3}$ are different states of one property (severity). Similarly, $P_{4},P_{5},P_{6}$ are different states of one property (mental), $P_{7},P_{8},P_{9}$ are different states of one property (risk), $P_{10},P_{11},P_{12}$ are different states of one property (action), Their simultaneous existence may refer to some concept of inconsistency in the rule base. The reachability graph has no loops means no circularity errors and finally having no transitions underlined, speaks for non-redundancy.
Regarding the principles’ coverage verification, by scanning the FPN, we can construct the following principles coverage vector:
$C=[\text{coverage(Autonomy)},\text{coverage(Beneficence)},\\ \text{coverage(Nonmaleficence)}]=[1,1,1]$
that proves that the model satisfies the principle coverage completeness. Regarding cross-principle consistency verification, in our case study we have the following principles pair [Autonomy, Nonmaleficence] is an incompatible pair. Looking at the reachability graph, we can see that no marking $M$ exists where a rule tagged Autonomy and a rule tagged Non-maleficence are simultaneously enabled with incompatible consequents. This indicates that no cross-principle conflict is detected.
Finally, no duplicate rules with identical semantic content were found. Example: $R_{1}$ and $R_{2}$ differ in both consequents and principles. $R_{4},R_{5},R_{6}$ have distinct risk preconditions and/or actions. Thus, no principle redundancy is detected.
@@
Figure 5: Reachability Graph of the FPN of the $PatientEDM^{+}$ model
@@ 5.2 Validation Process
To validate the created EDM model, we will need to construct validation referents. The validation referents provide a standard against which to compare our EDM models to determine the validity of these models for our needs. A validation referent gives the key information that a model must represent. Validation referents should be developed by experienced domain experts and expressed in the same language that is used to create the EDM model.
@@
Listing 3: Referent $V_{1}$ — Patient-Advocate / Autonomy-Focused
@@
7Principle ordering ($\Pi_{1}$): Autonomy $\succ$ Beneficence $\succ$ Nonmaleficence
@@
Listing 4: Referent $V_{2}$ — Clinician / Frontline Nurse (beneficence + nonmaleficence focus)
@@
7Principle ordering ($\Pi_{2}$): Nonmaleficence $\succ$ Beneficence $\succ$ Autonomy
@@
By applying the above steps on our $PatientEDM^{+}$ model, we found that: the property $LTconsequences$ which is an ethically relevant fact (mentioned by all three referent models) does not appear in the FPN of our $PatientEDM$ model, and also the rules C_R4, P_D1, P_D2, P_D3, C_D1, H_D1, H_D2, H_D3, H_D4 are missing. Therefore, we can conclude that our model is semantically incomplete with respect to the referent models. By adding these missing components we render our model semantically complete with respect to the validation referent models presented in listings 3, 4, 5 and discussed with the domain experts.
@@
Figure 6: FPN of the $PatientEDM^{+}$ Model
@@
This step involves checking for semantic incorrectness through running and reasoning over the FPN. We need to reason the FPN for given inputs and compare the results to their counterparts in the referent models to check if there exists any semantic incorrectness. The first step in this phase is to map the revised $PatientEDM^{+}$ model after static validation into FPN. Figure 6 shows the FPN of the revised model and Listing 6 shows its specifications.
Then, we will calculate action similarities between the $PatientEDM^{+}$ model output and the expected output by our three referent models. The similarities are calculated using the formulas from Section 3.4. Figure 7 shows the output of the $PatientEDM^{+}$ model together with the similarities to our three referents. Analyzing the current action similarity scores reveals validation failures: the Clinician’s $S_{A}=0.81$ falls short of their 0.85 threshold despite matching the expected tryAgainNow action, while Patient-Advocate ($S_{A}=0.49<0.75$) and Hospital-Board ($S_{A}=0.49<0.80$) reject tryAgainNow in favor of tryAgainLater under 63.1% crisp risk. To improve these similarities and boost validation rates, we can adjust certainty factors (CFs) in the FERD rules, for instance, increasing R5’s CF from 0.9 to 0.95 for stronger tryAgainNow emphasis (high risk), and elevating R6’s CF from 0.7 to 0.9 (or adding a new medium-risk rule like Dx: (risk(medium) $\wedge$ LTconsequences(medium) $\to$ tryAgainLater (CF: 0.9)) to better align with tryAgainLater expectations. These targeted modifications enhance system-referent coherence while preserving fuzzy semantic tolerances.
Figure 7: Output of the $PatientEDM^{+}$ model
Conversely, Figure 8 shows the output for the same scenario after the explained modification. After that, we will reason over the FPN applying the reasoning rules and comparing the results to the reasoning references ($RR1$ and $RR2$ in our case) from all three validation referents.
@@
Let us see RR1_V1RR1\_V1 in the Patient_Advocate validation referent in Listing 3. Using the FPN to reason, we can obtain the following reasoning results: $\alpha(Risk(high))=min(0.95,0.9)*0.9=0.9*0.9=0.81>0.80$. therefore, the validation rule of RR1_V1RR1\_V1 has passed through the validation. Let us see RR2_V1RR2\_V1 in the validation referent in Listing 3. Using the FPN to reason, we can obtain the following reasoning results: $\alpha(Action(Accept))=min(min(0.9,0.8)*0.8,0.85)=min(0.64,0.85)=0.64<0.70$. therefore, the validation rule of $RR2$ has not passed through the validation, which means that in the $PatientEDM^{+}$ model there are some semantic errors that need to be corrected. This error can be corrected/improved by, for example, increasing the CF of the rule $R1$ in the $PatientEDM^{+}$ model from 0.8 to 0.9 so $\alpha(Action(Accept))$ becomes $0.72>0.70$ and the model passes the second validation rule of the Patient_Advocate validation referent. Now we repeat the same with the reasoning rules of the other two referent models.
Figure 8: Output of the $PatientEDM^{+}$ model after modification
@@ 6 Discussion
6.1 Applicability to Reinforcement Learning and LLMs
While the proposed fEDM framework has been illustrated in a rule-based healthcare case study, an important question concerns its applicability to modern AI systems, particularly those based on reinforcement learning (RL) and large language models (LLMs). These paradigms are increasingly deployed in the safety-critical and human-facing domains, but often raise concerns about ethical reliability and transparency.
@@ 6.1 Applicability to Reinforcement Learning and LLMs
RL agents optimize policies with respect to a reward function, but this objective rarely captures the full range of ethical considerations relevant in human–AI interaction. Moreover, RL policies are opaque and can produce unintended behaviors when faced with out-of-distribution states. The fEDM+ framework can complement RL systems by functioning as an ethical oversight layer as detailed below:
The RL agent proposes an action $a$. This action, together with the current state’s ethically relevant factors (ERFs), is passed through the fEDM+ model. If the action produces an ethical risk level above a specified threshold, fEDM+ either blocks it or suggests alternatives consistent with ethical constraints. In this way, the RL agent retains adaptive learning capability, while fEDM+ enforces transparent, verifiable ethical boundaries.
@@
LLMs generate context-sensitive responses but are prone to producing ethically problematic or unsafe content. Directly constraining LLM generation is difficult due to the distributed nature of their representations. However, the fEDM+ framework can serve as a validation layer for LLM outputs: Candidate responses are evaluated with respect to fuzzy ethical rules (FERDs) instantiated from domain knowledge. The fEDM+ system performs risk assessment (via FERRs) and blocks or flags outputs exceeding acceptable risk thresholds. This provides a mechanism for embedding domain-specific ethical principles into LLM-based decision pipelines.
Advantages of Oversight Integration. By functioning as an external module, the fEDM+ framework offers three key advantages for modern AI: I) Transparency: Ethical constraints are explicitly encoded in fuzzy rules, unlike implicit reward shaping in RL or prompt engineering in LLMs. II) Verifiability: Rule bases can be formally checked for structural errors via FPN analysis, ensuring that oversight remains sound even as AI models evolve. III) Flexibility: Expert-defined rules can be supplemented with data-driven rule learning, enabling adaptation to dynamic domains while preserving ethical oversight.
Thus, while fEDM+ does not replace RL or LLM methods, it provides a practical and formally grounded mechanism for ensuring ethical reliability in their deployment. Future work will explore hybrid implementations where RL and LLM systems are monitored and constrained in real time by fEDM+ modules.